# Semantic Contextualization of Face Forgery: A New Definition, Dataset, and Detection Method

### v1

@@ Abstract
In recent years, deep learning has greatly streamlined the process of generating realistic fake face images. Aware of the dangers, researchers have developed various tools to spot these counterfeits. Yet none asked the fundamental question: What digital manipulations make a real photographic face image fake, while others do not? In this paper, we put face forgery in a semantic context and define that computational methods that alter semantic face attributes to exceed human discrimination thresholds are sources of face forgery. Guided by our new definition, we construct a large face forgery image dataset, where each image is associated with a set of labels organized in a hierarchical graph. Our dataset enables two new testing protocols to probe the generalization of face forgery detectors. Moreover, we propose a semantics-oriented face forgery detection method that captures label relations and prioritizes the primary task (i.e., real or fake face detection). We show that the proposed dataset successfully exposes the weaknesses of current detectors as the test set and consistently improves their generalizability as the training set. Additionally, we demonstrate the superiority of our semantics-oriented method over traditional binary and multi-class classification-based detectors.
@@ I Introduction
While several face forgery detection tools have been created, their generalizability to novel face manipulations remains limited. Early detectors rely primarily on three types of knowledge. The first is knowledge about statistical regularities of real face images [5, 6]. The second is knowledge about the digital photography pipeline, in which various visual artifacts may arise and can be characterized to expose forgeries [7, 8, 9, 10, 11]. The third is knowledge about the 3D world we are living in, particularly the physical laws that govern the interactions of light, optics, and objects [12, 13, 14, 15]. However, designing computational structures manually that exploit domain knowledge is a highly challenging task. Consequently, existing knowledge-driven methods are often tailored to specific forgery scenarios, resulting in limited generalizability. With the advent of deep learning, data-driven detectors [16, 17, 18, 19, 20, 21, 22, 23] have come to the forefront, whose effectiveness is attributed to the quality of training data and the formulation of face forgery detection.
@@
If we are unable to draw a boundary between real and fake manipulations, it is not possible to discuss the generalizability of face forgery detectors. Previous generation tests involve training detectors on some “deemed fake” manipulations (e.g., Deepfakes [1]) and measuring the performance on another set of “deemed fake” manipulations (e.g., Face2Face [26]). Clearly, this setting does not accurately reflect the complexities of real-world scenarios, which are far less constrained.
In this paper, we rethink face forgery from conceptual and computational perspectives. We first define face forgery in a semantic context:
@@
Here, the description “to exceed human discrimination thresholds” means that the alteration of semantic face attributes relative to the initial real photographic image is discernable to the human eye. Fig. 1 presents an example of diffusion autoencoders [27] on age manipulation. By moving the image latent along the age direction [^1] Such a direction can be identified as the weight vector of a linear classifier trained on latent codes of positive and negative images of the age attribute [27]., i.e., tuning the age parameter $d_{\mathrm{age}}$, we can generate a sequence of manipulated images, of which some are easily distinguishable from the original image [^2] The original real photographic face image is chosen from the FaceForensics++ dataset [28] for illustration purposes only. (e.g., $d_{\mathrm{age}}\leq-0.20$). In contrast, images with an age parameter greater than $-0.10$ are innocuous in real-world applications because they retain nearly all semantic face attributes.
@@
Figure 1: Illustrating of diffusion autoencoders [27] on age manipulation. By varying the age parameter $d_{\mathrm{age}}$, which controls the movement of the image latent along the age direction, we create a set of age-manipulated images, only a subset of which are considered fake according to our definition (e.g., those with $d_{\mathrm{age}}\leq-0.20$). A more negative number indicates an older age.
Guided by our definition, we create a new Face Forgery in the Semantic Context (FFSC) dataset, where each image in FFSC is associated with a set of semantic labels organized in a hierarchical acyclic graph, as shown in Fig. 2. We contextualize twelve popular face manipulations [27, 29, 30, 31, 32, 33, 20, 34, 35, 36, 37, 38] into five global face attributes (i.e., age, expression, gender, identity, and pose). Notably, a single manipulation method can modify multiple attributes, while multiple different methods can modify the same face attribute. In FFSC, the manipulation degree that exceeds human discrimination thresholds and the connectivity to the local face regions (i.e., eye, eyebrow, lip, mouth, nose, and skin) are determined using formal psychophysical testing. FFSC supports two fine-grained semantics-oriented testing protocols: 1) generalization to novel manipulation methods for the same face attribute, and 2) generalization to novel face attributes.
@@
  • a new face forgery detection method that is oriented to rely on face semantics.
@@ II-A Face Forgery Datasets
Four major face manipulation techniques are adopted in the construction of existing face forgery datasets. The first is face editing supplied by Adobe® Photoshop® (e.g., Face-Aware Liquify), which provides high-level semantic abstractions for face manipulations [41]. The second is face swapping based on autoencoders [1], which replaces the face from a target image/video with that from the source. The third is the application of conditional generative models [2, 42], based on generative adversarial networks (GANs) [43] and denoising diffusion models [2], for controllable face editing. The fourth is image-based face rendering [26], which estimates (morphable) 3D face models from input images/videos [44], followed by alignment and re-rendering. Representative face forgery datasets include UADFV [45], FaceForensics++ (FF++) [28], Celeb-DF [46], DeepFakeDetection (Google-DFD) [47], DeeperForensics-1.0 (DF-1.0) [48], DeepFake Detection Challenge (DFDC) [40], FFIW [49], ForgeryNet [24], and OW-DFA [25] with consistent improvements in dataset size, sample complexity, identity diversity, visual quality, task complexity, and ethical approval. Table I presents a detailed summary of these datasets, following the binary and $C$-way classification formulations. In stark contrast, the proposed FFSC dataset contextualizes face manipulations from the perspective of face semantics.
TABLE I: Summary of face forgery datasets
@@ II-B Face Forgery Detectors
Traditional forensics tools are designed to detect image forgeries by spotting statistical irregularities [5, 6], visual artifacts (such as demosaicking [7], chromatic aberration [8], vignetting [9], and noise [11]), and physical and geometric inconsistencies [12, 13, 14, 15] in the image. Face forgery detection has been significantly influenced by these methods as a subfield of forensic science. Generally, these knowledge-driven methods are limited by the expressiveness of handcrafted (and manipulation-specific) features.
With the advancements of deep learning, many methods learn to expose face forgeries from physiological signals, including eye blinking [16], head pose [45], pupil shape [50], corneal specularity [51], and behavioral patterns [52]. Pure data-driven approaches in the spatial [28, 17, 53, 54, 55, 20, 56, 57] and frequency [58, 59, 60] domain have also been proposed, coupling with advanced learning strategies, such as attention learning [21, 61, 62], adversarial learning [63], meta-learning [64], graph learning [65, 66], contrastive learning [67, 68], and multi-task learning [19, 25, 69, 70]. Empirically, data-driven methods tend to overfit training manipulations and struggle to generalize to novel manipulations.
Researchers have recently begun detecting fake faces by examining face semantics [71, 22]. Haliassos et al. [71] leveraged lipreading features, while Dong et al. [22] exploited face identity features. In this paper, we further explore this direction, and conduct a systematic reexamination of face forgery at the semantic level. The resulting SO-detectors can be loosely seen as a generalization of the above two methods, with significantly improved detection performance.
@@
Figure 3: Face manipulations adopted in the FFSC dataset. Each subfigure displays the original and manipulated face images on the left and right, respectively.
@@ III FFSC Dataset
In this section, we summarize our efforts to create the FFSC dataset, including data collection, face manipulation, and data annotation.
@@ III-A Data Collection
We start by collecting real photographic videos from two popular datasets, AVSpeech [72] and Celeb-DF [46] YouTube-real, which contain a diverse set of video clips of different ages, expressions, genders, identities, poses, ethnicities, and shooting conditions. We randomly select $700$ and $300$ high-resolution videos from the AVSpeech test set and Celeb-DF YouTube-real, respectively. As most face manipulation methods are image-based, we first uniformly sample $128$ frames from each video and detect the face regions in each frame by RetinaFace [73]. We only retain the largest face and extend it to $1.3^{2}$ times the area of the tight crop produced by the face detector, as suggested in [28]. We adjust each face image to a fixed size of $317\times 317$. Importantly, we exclude face images with low visual quality, closed eyes, extreme poses, and occlusions, which takes roughly a week by the first author and a research assistant at City University of Hong Kong. In total, we collect $63,344$ real face images, corresponding to $1,000$ face identities.
@@ III-B Face Manipulation
We construct the FFSC dataset by instantiating five global face attribute nodes: age, expression, gender, identity, and pose, through twelve face manipulation methods. We leverage eight face manipulation methods [27, 29, 30, 31, 32, 33, 20, 34] to build the main FFSC dataset, including the well-split training, validation, and test sets with the ratio of $7.8:1.1:1.1$. The independence of face identity is ensured during splitting. To create an additional test set, we reserve four face manipulation algorithms [38, 35, 36, 37]. This is designed to assess the generalizability of face forgery detectors when exposed to new manipulations that alter the same face attributes seen in the training set. To keep a data balance between real and fake face images in the main FFSC dataset, we randomly select one-eighth of real images for face manipulation, totaling $75,176$ fake images. As for the additional FFSC test set, we manipulate the real images from the validation and test sets of the main FFSC dataset to obtain $8,664$ fake images.
@@
Gender Manipulation. To flip the gender attribute, we adopt diffusion autoencoders [27] and StyleGAN2 distillation [31], i.e., converting males to females (see Fig. 3 (f)) and vice versa (see Fig. 3 (g)). Note that StyleGAN2 distillation is an image-to-image translation method that does not offer adjustable parameters to control the degree of manipulation.
Identity Manipulation. To alter the identity attribute, we follow the practice in existing face forgery datasets [28, 46, 40, 24], which swap two faces with different identities by SimSwap [32], FSGAN [33], and BlendFace [20]. The former two are data-driven methods based on deep learning, while the latter is a knowledge-driven algorithm. For each target face, we first search for the best source face by minimizing the Euclidean distance between detected face landmarks [20], while excluding faces with the same identity and different gender. Figs. 3 (h)-(j) show the visual examples, which involve no manipulation degree tuning.
Pose Manipulation. To alter the pose attribute, particularly the head posture in the horizontal plane, we adopt two methods: the first-order motion model [30] and the thin-plate spline motion model [34]. Similar to the former, the thin-plate spline motion model also requires a driving video as input to supply the horizontal head rotation movement, and allows more complex nonlinear motion transfer. The visual examples are shown in Figs. 3 (k) and (l).
As for the additional test set in FFSC, we employ four different face manipulation algorithms: HFGI [38] (for both age and expression attributes), StyleCLIP [35] (for the gender attribute), InfoSwap [36] (for the identity attribute), and FNeVR [37] (for the pose attribute). HFGI, StyleCLIP, and InfoSwap are GAN-based and only need a single face image as input, whereas FNeVR is based on motion transfer, which needs an additional driving video as input.
@@ III-C1 Specification of Manipulation Degree Parameters
For the face manipulation under consideration, we tune its manipulation degree parameter (if any) to generate a series of manipulated images. We then estimate human discrimination thresholds using a yes-no task. On each trial, three subjects are shown a real face image and a corresponding manipulated image (for one second and in randomized spatial order), and then asked to indicate whether the two images are perceptually different. This procedure is repeated for $300$ trials for each manipulation degree value and over $50$ image pairs, with ordering determined by a standard psychophysical staircase procedure [75]. The distribution of human responses, as a function of the manipulation degree parameter, is fitted using a Gaussian cumulative distribution function, and the human discrimination threshold is set to the parameter value such that the subjects can clearly perceive the visual differences between the two images $75\%$ of the time. Finally, we transfer the fitted degree parameter of each manipulation method (on the $50$ real images) to all remaining real images for fake face image generation.
@@
Figure 4: Human discrimination distributions of face attributes/regions. Zoom in for improved visibility.
@@ III-C2 Specification of Label Hierarchy
To link each fake image to a label hierarchy (refer to Fig. 2 (a)), it is essential to determine 1) whether altering one global face attribute impacts other non-targeted attributes and 2) which specific changes in local face regions result in the modification of the global face attribute (i.e., the connectivity between the global face attribute to the local face regions). Towards this goal, we invite a group of $12$ human subjects to participate in another formal psychophysical experiment. Similarly, a pair of real and manipulated images are shown to at least three subjects (for unlimited time and in randomized spatial order), who are asked to indicate whether the modification of non-targeted face attributes/regions is visually discriminable, each corresponding to a yes-no task. This process is repeated over $200$ image pairs for each face manipulation method, with human discrimination distributions shown in Fig. 4. We find that subjects are generally confident in perceiving non-targeted face attributes/regions as clearly discriminable (or indiscriminable) with a probability higher than $70\%$ (or lower than $30\%$), which is associated with a binary label of one (or zero). As for the non-targeted attributes where the probabilities of discrimination fall within the range of $30\%$ to $70\%$—for instance, the age attribute when adjusting for gender and identity—we refrain from assigning definitive binary labels. In such cases, we consider the underlying label of the non-targeted attribute to be unobserved. Additionally, the clear discrimination of targeted face attributes, each with a probability over $75\%$, verifies the specification of manipulation degree parameters in the preceding psychophysical experiment.
The results of the psychophysical experiment suggest a hierarchical acyclic graph instantiation to encode the semantic labels as shown in Fig. 2 (a). Starting from the root node face, we split it into five global face attribute nodes: age, expression, gender, identity, and pose, and connect them to a set of leaf nodes representing non-overlapping local face regions: eye, eyebrow, lip, mouth, nose, and skin. We share the label hierarchy, $\bm{y}\in\{0,1\}^{N}$, for images with the targeted face attribute manipulated by the same method. $y_{0}=1$ indicates that the test face image is fake, while $y_{0}=0$ signifies that it is real. Likewise, $y_{i}=1$, for $i>1$, indicates the $i$-th face attribute/region is fake, and $y_{i}=0$ otherwise.

### v3

@@ Abstract
In recent years, deep learning has greatly streamlined the process of manipulating photographic face images. Aware of the potential dangers, researchers have developed various tools to spot these counterfeits. Yet, none asks the fundamental question: What digital manipulations make a real photographic face image fake, while others do not? In this paper, we put face forgery in a semantic context and define that computational methods that alter semantic face attributes to exceed human discrimination thresholds are sources of face forgery. Following our definition, we construct a large face forgery image dataset, where each image is associated with a set of labels organized in a hierarchical graph. Our dataset enables two new testing protocols to probe the generalizability of face forgery detectors. Moreover, we propose a semantics-oriented face forgery detection method that captures label relations and prioritizes the primary task (i.e., real or fake face detection). We show that the proposed dataset successfully exposes the weaknesses of current detectors as the test set and consistently improves their generalizability as the training set. Additionally, we demonstrate the superiority of our semantics-oriented method over traditional binary and multi-class classification-based detectors.
@@ I Introduction
While several face forgery detection tools have been created, their generalizability to novel face manipulations remains limited. Early detectors rely primarily on three types of knowledge. The first is knowledge about statistical regularities of real face images [5, 6]. The second is knowledge about the digital photography pipeline, in which various visual artifacts may be characterized to expose forgeries [7, 8, 9, 10, 11]. The third is knowledge about the 3D world we are living in, particularly the physical laws that govern the interactions of light, optics, and objects [12, 13, 14, 15]. However, designing computational structures manually to exploit domain knowledge is a highly challenging task. Consequently, existing knowledge-driven methods are often tailored to specific forgery scenarios, resulting in limited generalizability. With the advent of deep learning, data-driven detectors [16, 17, 18, 19, 20, 21, 22, 23] have come to the forefront, whose effectiveness is attributed to the quality of training data and the formulation of face forgery detection.
@@
If we are unable to draw a boundary between real and fake manipulations, it is not possible to discuss the generalizability of face forgery detectors. Previous generalization tests involve training detectors on some “deemed fake” manipulations (e.g., Deepfakes [1]) and measuring performance on another set of “deemed fake” manipulations (e.g., Face2Face [26]). Clearly, this setting does not accurately reflect the complexities of real-world scenarios.
In this paper, we rethink face forgery from both conceptual and computational perspectives. We first define face forgery in a semantic context:
@@
Here, the description “to exceed human discrimination thresholds” means that the alteration of semantic face attributes relative to the original real photographic image is discernable to the human eye. Fig. 1 presents an example of diffusion autoencoders [27] on age manipulation. By moving the image latent along the age direction [^1] Such a direction can be identified as the weight vector of a linear classifier trained on latent codes of positive and negative images of the age attribute [27]., i.e., tuning the age parameter $d_{\mathrm{age}}$, we can generate a sequence of manipulated images, of which some are easily distinguishable from the original image [^2] The original real photographic face image is chosen from the FaceForensics++ dataset [28] for illustration purposes only. (e.g., $d_{\mathrm{age}}\leq-0.20$). In contrast, images with the age parameter greater than $-0.10$ are innocuous in real-world applications because they retain nearly all semantic face attributes.
@@
Figure 1: Illustration of diffusion autoencoders [27] on age manipulation. By varying the age parameter $d_{\mathrm{age}}$, which controls the movement of the image latent along the age direction, we create a set of age-manipulated images, only a subset of which are considered fake according to our definition (e.g., those with $d_{\mathrm{age}}\leq-0.20$). A more negative number indicates an older age.
Guided by our definition, we create a new Face Forgery in the Semantic Context (FFSC) dataset, in which each image is associated with a set of semantic labels organized in a hierarchical acyclic graph, as shown in Fig. 2. We contextualize twelve popular face manipulations [27, 29, 30, 31, 32, 33, 20, 34, 35, 36, 37, 38] into five global face attributes (i.e., age, expression, gender, identity, and pose). Notably, a single manipulation method can modify multiple attributes, while multiple different methods can modify the same face attribute. In FFSC, the manipulation degree that exceeds human discrimination thresholds and the connectivity to local face regions (i.e., eye, eyebrow, lip, mouth, nose, and skin) are determined using formal psychophysical testing. FFSC supports two fine-grained semantics-oriented testing protocols: 1) generalization to novel manipulation methods for the same face attribute, and 2) generalization to novel face attributes.
@@
  • a new face forgery detection method that is oriented to relying on face semantics.
@@ II-A Face Forgery Datasets
Four major face manipulation techniques are adopted in the construction of existing face forgery datasets. The first is face editing supplied by Adobe® Photoshop® (e.g., Face-Aware Liquify), which provides high-level semantic abstractions for face manipulations [41]. The second is face swapping based on autoencoders [1], which replaces the face from a target image/video with that from the source. The third is the application of conditional generative models [2, 42], such as generative adversarial networks (GANs) [43] and denoising diffusion models [2] for controllable face editing. The fourth is image-based face rendering [26], which estimates (morphable) 3D face models from input images/videos [44], followed by alignment and re-rendering. Representative face forgery datasets include UADFV [45], FaceForensics++ (FF++) [28], Celeb-DF [46], DeepFakeDetection (Google-DFD) [47], DeeperForensics-1.0 (DF-1.0) [48], DeepFake Detection Challenge (DFDC) [40], FFIW [49], ForgeryNet [24], KoDF [50], DF-Platter [51], and OW-DFA [25] with consistent improvements in dataset size (ranging from hundreds [45] to tens of thousands [24]), sample complexity (shifting from single-person to multi-person face forgery [49, 51]), identity diversity (growing from tens [45, 46] to thousands [24]), visual quality (improving from noticeable artifacts to photorealistic outputs), task complexity (evolving from binary classification [28], $C$-way classification [25, 24, 52] to forgery localization [49, 24]), and ethical approval. Table I presents a summary of these datasets, following the binary and $C$-way classification formulations. In contrast, the proposed FFSC dataset contextualizes face manipulations from the perspective of face semantics.
TABLE I: Summary of existing face forgery datasets
@@ II-B Face Forgery Detectors
Traditional forensics tools are designed to detect image forgeries by spotting statistical irregularities [5, 6], visual artifacts (such as demosaicking [7], chromatic aberration [8], vignetting [9], and noise [11]), and physical and geometric inconsistencies [12, 13, 14, 15]. Face forgery detection has been significantly influenced by these methods as a subfield of forensic science. Generally, these knowledge-driven methods are limited by the expressiveness of handcrafted (and manipulation-specific) features.
With the advancements of deep learning, many methods learn to expose face forgery from physiological signals, including eye blinking [16], head pose [45], pupil shape [53], corneal specularity [54], and behavioral patterns [55]. Pure data-driven approaches in the spatial [28, 17, 56, 57, 58, 20, 59, 60] and frequency [61, 62, 63] domain have also been proposed, coupling with advanced learning strategies, such as attention learning [21, 64, 65], adversarial learning [66], meta-learning [67], graph learning [68, 69], contrastive learning [70, 71, 72, 73], and multi-task learning [19, 25, 74, 75, 76]. Empirically, data-driven methods tend to overfit training manipulations and struggle to generalize to novel manipulations.
Only until recently have researchers begun detecting fake faces through analysis of face semantics [77, 22]. Haliassos et al. [77] leveraged lipreading features, while Dong et al. [22] exploited face identity features. In this paper, we further explore this direction, and conduct a systematic reexamination of face forgery at the semantic level. The resulting SO-detectors can be loosely seen as a generalization of the above two methods, with significantly improved detection performance.
@@
Figure 3: Face manipulation techniques adopted in FFSC. Each subfigure displays the original and manipulated face images on the left and right, respectively.
@@ III FFSC Dataset
In this section, we introduce the FFSC dataset, including data collection, face manipulation, and data annotation.
@@ III-A Data Collection
We start by collecting real photographic videos from two popular datasets, AVSpeech [78] and Celeb-DF YouTube-real [46], which contain a diverse set of video clips of different ages, expressions, genders, identities, poses, ethnicities, and shooting conditions. We randomly select $700$ and $300$ high-resolution videos from the AVSpeech test set and Celeb-DF YouTube-real [^3] Celeb-DF Youtube-real is an additional subset for extended use, and does not overlap with the official Celeb-DF dataset., respectively. As most face manipulation methods are image-based, we first uniformly sample $128$ frames from each video and detect the face regions in each frame by RetinaFace [79]. We only retain the largest face and extend it to $1.3^{2}$ times the area of the tight crop produced by the face detector [28]. We adjust each face image to a fixed size of $317\times 317$. Importantly, we exclude face images with low visual quality, closed eyes, extreme poses, and occlusions. This meticulous screening process, conducted by the first author and a research assistant at the City University of Hong Kong, takes approximately one week to complete. In total, we collect $63,344$ real face images, corresponding to $1,000$ face identities.
@@ III-B Face Manipulation
We construct the FFSC dataset by instantiating five global face attribute nodes: age, expression, gender, identity, and pose, through twelve face manipulation methods. We leverage eight methods [27, 29, 30, 31, 32, 33, 20, 34] to build the main FFSC dataset, including the well-split training, validation, and test sets with a ratio of $7.8:1.1:1.1$. The independence of face identity is ensured during splitting. We reserve four face manipulation algorithms [38, 35, 36, 37] to create an additional test set. This is designed to assess the generalizability of face forgery detectors when exposed to new manipulations that alter the same face attributes in the training set. To balance between real and fake face images in the main FFSC dataset, we randomly select one-eighth of real images for face manipulation, totaling $75,176$ fake images. As for the additional FFSC test set, we manipulate the real images from the validation and test sets of the main FFSC dataset to obtain $8,664$ fake images.
@@
Gender Manipulation. To flip the gender attribute, we adopt diffusion autoencoders [27] and StyleGAN2 distillation [31], converting males to females (see Fig. 3(f)) and vice versa (see Fig. 3(g)). Note that StyleGAN2 distillation is an image-to-image translation method that does not offer adjustable parameters to control the degree of manipulation.
Identity Manipulation. To alter the identity attribute, we follow the practice in existing face forgery datasets [28, 46, 40, 24], which swaps two faces with different identities by data-driven SimSwap [32] and FSGAN [33], and knowledge-driven BlendFace [20]. For each target face, we first search for the most suitable source face by minimizing the Euclidean distance between detected face landmarks [20], while excluding faces with the same identity and different gender. Figs. 3(h)-(j) show the visual examples, which involve no manipulation degree tuning.
Pose Manipulation. To alter the pose attribute, particularly the head posture in the horizontal plane, we adopt two methods: the first-order motion model [30] and the thin-plate spline motion model [34]. Similar to the former, the spline motion model also requires a driving video as input and allows more complex nonlinear motion transfer. The visual examples are shown in Figs. 3(k) and (l).
As for the additional test set in FFSC, we employ four different face manipulation algorithms: HFGI [38] (for both age and expression attributes), StyleCLIP [35] (for the gender attribute), InfoSwap [36] (for the identity attribute), and FNeVR [37] (for the pose attribute). HFGI, StyleCLIP, and InfoSwap are GAN-based and only need a single face image as input, whereas FNeVR relies on motion transfer, requiring an additional driving video as input.
@@ III-C1 Specification of Manipulation Degree Parameters
For the face manipulation under consideration, we tune its manipulation degree parameter (if any) to generate a series of manipulated images. We then estimate human discrimination thresholds using a standardized psychophysical procedure [81, 82]. On each trial, three subjects are shown a real face image as the reference and a corresponding manipulated image as the stimulus (for one second and in randomized spatial order), and then asked to indicate whether the two images are perceptually different. This procedure is repeated for $300$ trials for each manipulation degree value and over $50$ image pairs, with ordering determined by a standard staircase method [81, 82]. The distribution of human responses, as a function of the manipulation degree parameter, is fitted using a Gaussian cumulative distribution function, and the human discrimination threshold is set to the parameter value such that the subjects can clearly perceive the visual differences between the two images $75\%$ of the time [82]. Finally, we transfer the fitted degree parameter of each manipulation method (on the $50$ real images) to all remaining real images.
@@
Figure 4: Human discrimination distributions of face attributes and regions. Zoom in for improved visibility.
@@ III-C2 Specification of Label Hierarchy
To pair each fake image with a label hierarchy (refer to Fig. 2(a)), it is essential to determine 1) whether altering one global face attribute impacts other non-targeted attributes and 2) which specific changes in local face regions result in the modification of the global face attribute (i.e., the connectivity between the global face attribute to local face regions). Toward this goal, we invite a group of $12$ human subjects to participate in another formal psychophysical experiment. Similarly, a pair of real and manipulated images are shown to at least three subjects (for unlimited time and in randomized spatial order) to indicate whether the modification of non-targeted face attributes (and regions) is visually discriminable, each corresponding to a yes-no task. This process is repeated over $200$ image pairs for each face manipulation method, with human discrimination distributions shown in Fig. 4. We find that subjects are generally confident in perceiving the changes in non-targeted face attributes as clearly discriminable (or indiscriminable) with probabilities higher than $70\%$ (or lower than $30\%$), corresponding to a binary label of one (or zero). As for the non-targeted attributes where the probabilities of discrimination fall within the range of $30\%$ to $70\%$—for instance, the age attribute when adjusting for gender and identity—we refrain from assigning definitive binary labels. In such cases, we consider the underlying label of the non-targeted attribute to be unobserved. Additionally, the clear discrimination of targeted face attributes, each with a probability over $75\%$, verifies the specification of manipulation degree parameters in the preceding psychophysical experiment.
The psychophysical results suggest a hierarchical acyclic graph instantiation to encode the semantic labels as shown in Fig. 2(a). Starting from the root node face, we split it into five global face attribute nodes: age, expression, gender, identity, and pose, and connect them to a set of leaf nodes representing non-overlapping local face regions: eye, eyebrow, lip, mouth, nose, and skin. We share the label hierarchy $\bm{y}\in\{0,1\}^{N}$, for images with the targeted face attribute manipulated by the same method. $y_{0}=1$ indicates that the test face image is fake, and vice versa. Likewise, $y_{i}=1$, for $i>1$, indicates the $i$-th face attribute or region is fake, and $y_{i}=0$ otherwise.
The proposed SO-contextualization for face manipulations allows a single manipulation to alter multiple face attributes (e.g., age and expression by StyleRes [29]). Meanwhile, different manipulations can affect the same attribute (e.g., expression by FOMM [30] and StyleRes). This contextualization discourages the detectors from relying on “shortcuts” as observed in [52], and instead facilitates the learning of more generalizable features.